### v1

# DAOcc: 3D Object Detection Assisted Multi-Sensor Fusion for 3D Occupancy Prediction

Abstract
Multi-sensor fusion significantly enhances the accuracy and robustness of 3D semantic occupancy prediction, which is crucial for autonomous driving and robotics. However, existing approaches depend on large image resolutions and complex networks to achieve top performance, hindering their application in practical scenarios. Additionally, most multi-sensor fusion approaches focus on improving fusion features while overlooking the exploration of supervision strategies for these features. To this end, we propose DAOcc, a novel multi-sensor fusion occupancy network that leverages 3D object detection supervision to assist in achieving superior performance, while using a deployment-friendly image feature extraction network and practical input image resolution. Furthermore, we introduce a BEV View Range Extension strategy to mitigate the adverse effects of reduced image resolution. As a result, our approach achieves new state-of-the-art results on the Occ3D-nuScenes and SurroundOcc datasets, using ResNet50 and a 256x704 input image resolution. Code will be made available at https://github.com/AlphaPlusTT/DAOcc.
1 Introduction
Figure 1: Unlike existing multi-modal methods [46, 33, 58, 40, 34], which rely on complex combinations of loss functions, including several or all of the focal loss $\mathcal{L}_{focal}$ [38], the scene-class affinity loss $\mathcal{L}_{scal}^{geo}$ and $\mathcal{L}_{scal}^{sem}$ [4], the lovasz-softmax loss $\mathcal{L}_{ls}$ [2], and the depth loss $\mathcal{L}_{d}$ [21], to achieve optimal performance, our approach only requires a single cross-entropy loss $\mathcal{L}_{ce}$ for the occupancy prediction task.
3D semantic occupancy prediction (occ) is a critical task in autonomous driving [51, 4, 47] and robotic systems [42, 44, 14, 45], where accurately understanding the environment is essential for safe and efficient navigation. Reliable occupancy prediction requires not just accurate spatial data but also a comprehensive understanding of the environment’s context. Achieving this necessitates the integration of data from multiple sensors. LiDAR provides precise 3D spatial information for obstacle detection, while cameras capture visual details like color and texture for a deeper understanding of the scene. By combining these complementary data sources, the accuracy and robustness of occupancy prediction are significantly enhanced.
In existing works on multi-modal [57, 33, 58, 49, 40, 46, 34] or image-based [4, 48, 41, 12, 47, 23, 54, 39, 50] occ tasks, achieving superior performance often involves using extremely high image resolutions and complex image feature extraction networks (see Table 1 and Table 3), such as using 900×1600 resolution input image and ResNet101 [11] equipped with DCN [7, 62]. However, this approach significantly limits the deployment of top-performing occ models on edge devices due to the high computational demands. Compared to image, point cloud is much sparser. For example, in the training set of the nuScenes dataset, the maximum number of points in a single frame point cloud is only 34,880, which is equivalent to just 2.4% of the number of pixels in a 900×1600 resolution image. Therefore, how to effectively leverage point cloud data within a multi-modal occ framework remains to be further explored.
Moreover, we observe that most works [57, 33, 58, 49, 40, 46] on multi-modal occ primarily focuses on obtaining more effective fusion features, with insufficient exploration into the forms of supervision for these fused features. Although CO-Occ [34] introduces regularization based on implicit volumetric rendering to supervise fused features, it only utilizes distance ground truth from the original point cloud data, failing to fully exploit the geometric and structural information inherent in point cloud features. In contrast, point cloud-based 3D object detectors [56, 6, 31] effectively leverage this information, achieving significantly better performance in 3D object detection tasks compared to image-based detectors [26, 16, 27, 28]. This observation suggests a new research direction: how to better exploit the unique strengths of point cloud data in multi-modal occ tasks.
Figure 2: Overview of our proposed DAOcc. We first introduce the BVRE strategy to enrich the spatial contextual semantic information by broadening the perceptual range from the BEV perspective. Then feature extraction is performed through the multi-modal occupancy prediction network illustrated in the above figure. Furthermore, to fully leverage the inherent geometric structures within point cloud features, we incorporate 3D object detection as auxiliary supervision.
Based on these observations, we propose DAOcc, a novel multi-modal occ framework that leverages 3D object detection to assist in achieving superior performance while using a deployment-friendly image feature extraction network and practical input image resolution.
In constructing the baseline network structure of DAOcc, we adopt the most direct and straightforward approach: first, we extract features from images and point clouds, respectively. Given that depth estimation from monocular images is an ill-posed problem [13, 20] and that deformable attention modules are overly complex [10], we employ a simple method, similar to Harley et al. [9], to transform image features from 2D space to 3D volumetric space. Specifically, we project a set of predefined 3D points onto the 2D image feature plane and use bilinear interpolation to sample the corresponding 2D image features for these 3D points. Next, we adopt the same simplest fusion strategy as BEVFusion. [30], concatenating image and point cloud features, then performing a 2D convolution to fuse them, leading to a unified BEV features. Finally, we apply a fully convolutional BEV encoder with a residual structure to further fuse the unified BEV features, and subsequently employ the Channel-to-Height operation [54] to reshape the height of the unified BEV features.
To fully exploit the geometric and structural information inherent in point cloud features, we augment the baseline model’s unified BEV features with 3D object detection supervision, thereby enhancing the discriminability of the unified BEV features. This renders the unified BEV features more sensitive to object boundaries and capable of perceiving relationships between internal object structures. Additionally, given the sparsity of point clouds, we extend the processing range of the point clouds and employ sparse convolutions [52] to mitigate the computational overhead introduced by this extension. We refer to this approach as BVRE (BEV View Range Extension). BVRE provides a larger BEV field of view, offering more contextual information and mitigating the adverse effects of reduced image resolution. It is worth noting that the 3D object detection supervision is only used as an auxiliary branch during training and can be removed during inference for the occ task.
As a result, the proposed DAOcc establishes the new state-of-the-art performance on the Occ3D-nuScenes and SurroundOcc benchmarks, while using ResNet50 and a 256×704 input image resolution. Specifically, on the Occ3D-nuScenes validation set, DAOcc achieves 53.82 mIoU when using the camera mask during training and 48.2 RayIoU without using the camera mask. Additionally, on the SurroundOcc validation set, DAOcc achieves 45.0 IoU and 30.5 mIoU.
In conclusion, our contributions are summarized as follows:
  • We design a simple yet efficient multi-modal 3D semantic occupancy prediction baseline, eliminating the need for complex deformable attention modules as well as image depth estimation during feature fusion.
  • We propose DAOcc, a novel multi-modal occupancy prediction framework that leverages 3D object detection to assist in achieving superior performance while using a deployment-friendly image feature extraction network and input image resolution.
  • We introduce a BEV View Range Extension strategy, which provides a larger BEV field of view, offering more contextual information and mitigating the adverse effects of reduced image resolution.
  • We establish the new state-of-the-art performance on Occ3D-nuScenes dataset and SurroundOcc dataset, while using ResNet50 and a 256×704 input image resolution.
2 Related Work
3D Occupancy Prediction: 3D occupancy prediction aims to map all occupied voxels in the environment and assign semantic labels, thus providing more fine-grained perception results.
Camera-based 3D occupancy prediction has gained significant attention due to its cost-effectiveness. MonoScene [4] is the first work to infer the dense occupancy and semantics of outdoor and indoor scenes from a single RGB image. Since full perception of the surrounding environment is crucial for making accurate decisions in autonomous driving and robotics, most of the recent work is based on surround view image input. TPVFormer [15] proposes an efficient tri-perspective view (TPV) representation that combines BEV with two additional perpendicular planes to provide a 3D perception result with multi-view image inputs. Given the insufficient fine-grained semantic information of the TPV representation, OccFormer [59] leverages dense 3D features and proposes a dual-path transformer-based occupancy network. However, both TPVFormer and OccFormer use sparse LiDAR points as supervision, resulting in sparse occupancy predictions. To obtain dense occupancy prediction, OpenOccupancy [46], SurroundOcc [48] and Occ3D [41] developed methods for generating dense occupancy labels and established benchmarks on their respective proposed datasets. It is worth noting that SurroundOcc does not introduce manual annotation. The emergence of these benchmarks further promoted the development of 3D occupancy prediction. PanoOcc [47] adopts voxel query and a coarse-to-fine approach to learn a unified occupancy representation. FlashOcc [54] uses BEV features with 2D convolution to achieve efficient feature extraction and proposes the channel-to-height transformation to lift the output of BEV to 3D space. SparseOcc [29] leverages a sparse voxel decoder to reconstruct the sparse geometry of a scene and uses sparse queries to predict masks and labels. OSP [39] treats the scene as a collection of points and uses flexible sampling to allow the model to pay special attention to certain areas. However, all camera-based methods inevitably need to cope with the challenges of adverse lighting and weather conditions, and many works inevitably lack robustness due to the ill-posed monocular depth estimation problem [13, 20].
While camera-based occupancy prediction has demonstrated promising results, multi-modal approaches offer greater reliability and robustness, making them indispensable for practical applications in autonomous driving and robotics. Due to the fact that cameras are susceptible to changing lighting and weather conditions, OccFusion [33] enhances the accuracy and robustness of the occupancy network by integrating features from LiDAR and radar. In concurrent work, OccFusion [58] projects the preprocessed denser and more uniform point cloud onto the image plane to establish a mapping relationship, and performs deformable attention [63] to fuse the corresponding features. Although OccFusion [58] avoids depth estimation, using deformable attention incurs a greater computational burden. Hydra [49] extends FB-BEV [24] to the camera-radar fusion domain and improves the accuracy of depth estimation with the help of radar features. Co-Occ [34] employs K-nearest neighbor search in selecting neighboring camera features to enhance the corresponding LiDAR features and proposes a regularization based on implicit volume rendering. However, this regularization only utilizes the distance ground truth of the point cloud and does not leverage its intrinsic geometry information. EFFOcc [40] proposes an efficient and lightweight multi-modal 3D occupancy network, but it significantly relies on the 3D detection pretraining of the LiDAR branch for optimal results. This dependence, however, limits the flexibility of the network structure design. Specifically, to maximize the benefits of the 3D detection pretraining, EFfOcc has to adjust its network architecture to align more closely with well-established LiDAR detection networks. In addition, the inherently ill-posed nature of the monocular depth estimation problem inevitably renders the image branch of EFFOcc less robust [13]. In this work, we introduce a simple yet efficient multi-modal occupancy prediction network, eliminating the need for complex deformable attention [63, 22] as well as image depth estimation during feature fusion, and does not require 3D object detection pretraining. Furthermore, to fully exploit the inherent geometry information in point cloud features, we add supervision of 3D object detection as an auxiliary branch on the fused features.
Multi-Modal 3D Object Detection: Recent multi-modal 3D object detection methods [30, 18, 19, 5, 43] mainly focus on learning effective BEV feature representations. TransFusion [1] proposes a two-stage transformer-decoder based detection head and applies cross attention to obtain image features for each object query. BEVFusion [30] proposes an efficient and generic multi-task multi-sensor fusion framework which unifies multi-modal features in the shared bird’s-eye view (BEV) representation space, and introduces a specialized kernel to speed up the BEV pooling operation. In concurrent work, to improve the robustness of the LiDAR-camera fusion framework to sensor failures, such as missing LiDAR sensor input, BEVFusion decomposes the LiDAR-camera fusion into two streams that can independently output perception results, and performs feature fusion after the two streams. DAL [13] follows the concept of ‘Detecting As Labeling’ and decouples the fused features during classification and regression. Specifically, it employs the fused features for classification, whereas it utilizes the point cloud features exclusively for regression. In this work, we incorporate the simple yet effective feature fusion approach of BEVFusion and introduce 3D object detection as an auxiliary branch during training.
3 Proposed Method
3.1 Overall Framework
Our goal is to fully exploit the advantages of point cloud features for multi-modal occupancy prediction. Previous multi-modal studies have not fully exploited this and can only achieve superior performance through more complex image feature extraction networks and larger input image resolutions. The overall framework of our proposed DAOcc is illustrated in Figure 2. DAOcc takes surround images and their corresponding time-synchronized point cloud as input, and obtains the features of the image and point cloud through the Camera Encoder and LiDAR Encoder, respectively. The 2D image features are transformed into the 3D voxel space by projection and sampling. Subsequently, the image and point cloud features in the 3D space are compressed along the height dimension to generate the corresponding BEV features. A simple 2D convolution is then applied for feature fusion, and the fully convolutional BEV Encoder encodes these fused features to obtain the final BEV representations. Finally, the occupancy head uses the Channel-to-Height operation [54] to restore the height of the BEV representations, resulting in final 3D voxel space representations that can be utilized for occupancy prediction. These modules collectively form the basic network architecture of DAOcc, which will be elaborated in Section 3.2.
On the foundation of the basic network, we introduce the BVRE strategy (see Section 3.3) to compensate for the information loss that arises from the reduction in image resolution. This strategy aims to enrich the spatial contextual semantic information by broadening the perceptual range from the Bird’s Eye View (BEV) perspective. Furthermore, to fully leverage the inherent geometric structures within point cloud features, we incorporate 3D object detection as auxiliary supervision (see Section 3.4). This auxiliary supervision not only enhances the discriminability of the fused features but also leads to a very concise overall training loss (see Section 3.5) for our proposed framework.
3.2 Basic Network
LiDAR Encoder. The method of embedding raw LiDAR points into 3D voxelized features is consistent with BEVFusion [30]. We first voxelize the point cloud, retaining a maximum of 10 points per voxel, resulting in a 3D voxel grid of size $D\times H\times W$.The feature representation for each voxel is obtained by averaging the features of all points within it. Next, we apply 3D sparse convolutions [52] to encode these voxel features, generating spatially compressed LiDAR voxel features $F_{l}\in\mathbb{R}^{C\times\frac{D}{16}\times\frac{H}{8}\times\frac{W}{8}}$, where $C$ represents the feature dimensions.
Camera Encoder. For image feature extraction, taking surround images as input, we first use ResNet50 [11] as the backbone to extract multi-scale features, denoted as $F_{ms}=\{F_{ms\times\frac{1}{8}},F_{\frac{1}{16}},F_{\frac{1}{32}}\}$, where $F_{\frac{1}{x}}$ represent features extracted after $x\times$ downsampling. Then, we employ the Feature Pyramid Network (FPN) [25] as the neck to aggregate these multi-scale features. The output feature map $F_{c_{p}}$ has a shape of $N_{p}\times C_{p}\times\frac{H_{p}}{8}\times\frac{W_{p}}{8}$, where $H_{p}$ and $W_{p}$ represent the input resolution of the image, and $C_{p}$ and $N_{p}$ indicate the number of channels and the number of surround images, respectively.
Projection and Interpolation. For image-related occupancy prediction, a key step is to transform image features from 2D image plane to 3D volume space. Most existing methods use monocular depth estimation [46, 40, 12, 54] or deformable attention [48, 41, 47, 58]. However, monocular depth estimation is inherently an ill-posed problem [13, 20], while deformable attention imposes a significant computational burden [10]. Given these limitations, we use a simple yet effective projection and sample approach similar to Harley et al. [9]. Concretely, we first predefine a 3D voxel grid with a shape of $Z\times\frac{H}{8}\times\frac{W}{8}$, where $Z$ represents the number of voxels along the $z$-axis. The center point of each voxel is then projected onto the image feature plane, and only points that fall within both the image feature plane and the camera’s field of view are retained. Next, the sub-pixel projection positions of the retained points are bilinearly interpolated to generate the image features corresponding to each voxel. For voxels located in the overlapping viewing regions of the surround cameras, we average the image features from the two corresponding cameras to obtain the final feature for each voxel. The output camera voxel features can be denoted as $F_{c}\in\mathbb{R}^{C\times Z\times\frac{H}{8}\times\frac{W}{8}}$.
BEV Encoder. Given the fused feature $F_{f}$, we further refine $F_{f}$ by passing it through three blocks of ResNet18 [11], resulting in two feature maps, $F_{f0}$ and $F_{f2}$, extracted at two scales from the first and last blocks respectively. Then, similar to FPN [25], we apply bilinear upsampling to $F_{f2}$ and concatenate it with $F_{f0}$ along the feature dimension. Finally, we fuse the features of different scales using a convolution block. The output refined BEV features can be denoted as $F_{r}\in\mathbb{R}^{C_{r}\times\frac{H}{8}\times\frac{W}{8}}$.
3.3 BVRE
Considering the sparsity of point clouds, the increase in computational cost resulting from expanding the processing range of point clouds is very small compared to images. Therefore, we extend the point cloud range to provide more 3D spatial context, compensating for the information loss caused by lower image resolution. However, arbitrarily setting the $XY$ range of the point cloud or voxel resolution may result in misalignment between the BEV features and the occupancy ground truth. Assuming that the voxel resolution of the occupancy grid is $res_{o}$, the expanded point cloud data in BEV covers a rectangular region with $x$-coordinates ranging from $-x$ to $x$ and $y$-coordinates ranging from $-y$ to $y$, and has a voxel resolution of $res_{p}$. Thus, $x$ and $y$ must be integer multiples of $res_{o}$, and $res_{o}$ must also be an integer multiple of $res_{p}$. Otherwise, the spatial resolution along the $z$-axis for each feature vector in the BEV representation will not be equal to $res_{o}$. To avoid complex manual design, we adopted coordinate transformation and interpolation, as illustrated in the purple part of Figure 2, which can be formulated as follows:
$F_{occ}={\rm GridSample}(F_{r},{\rm Norm}(T_{o2p}\times P_{o}))$ (1)
where $P_{o}$ is a set of predefined points in the $XY$ plane of the coordinate system of the occupancy annotation, each point is located at the center of an occupancy grid in the $XY$ plane. $T_{o2p}$ is the transformation matrix from occupancy coordinates to Lidar coordinates. The function of ${\rm Norm}$ is to scale the coordinate values to a range from -1 to 1.
3.4 Detachable Auxiliary Head
To fully exploit the inherent geometry information in point clouds, we add a separable auxiliary detection network to the fused features, which will enable the features to benefit more from the geometry information. Meanwhile, the auxiliary detection task is related to the occupancy prediction task, thus providing multiple regularization effects during optimization. For simplicity, we use the one-stage CenterHead introduced in CenterPoint [53] as the auxiliary detection head. Given the refined BEV feature $F_{r}$, we utilize two convolutional layers to generate a keypoint heatmap $HM={p_{xy}}$ from the BEV perspective, and apply a Gaussian kernel [17, 60] to map the center points of all ground truth 3D boxes onto a target heatmap $T$. The training objective is formulated as a focal loss [38] based on Gaussian heatmaps:
$\mathcal{L}_{cls}=\frac{-1}{N}\sum_{ij}\begin{cases}(1-p_{ij})^{\alpha}\log(p_{ij})&\!\text{if}\ y_{ij}=1\\ (1-y_{ij})^{\beta}(p_{ij})^{\alpha}\log(1-p_{ij})&\text{otherwise}\end{cases}$ (2)
where $p_{ij}$ and $y_{ij}$ represent the predicted score and ground truth of the heatmap at location $(i,j)$, respectively. $N$ is the number of objects in a point cloud, and $\alpha$ and $\beta$ are hyper-parameters of the focal loss [38]. Similarly, for the offset of the center point, as well as the height-above-ground, 3D dimensions and yaw of the 3D bounding box, we use separate convolutional layers to predict each of these parameters, and then apply the L1 loss for supervision:
$\mathcal{L}_{loc}=\frac{1}{N}\sum_{k=1}^{N}\left|\hat{s}_{k}-s_{k}\right|$ (3)
where $\hat{s}_{k}$ and $s_{k}$ represent the predicted value and the ground truth respectively. The above two tasks enable the network to perceive object boundaries more precisely, which in turn facilitates achieving more precise occupancy predictions. The total loss for the auxiliary detection head can be denoted as:
$\mathcal{L}_{det}=\mathcal{L}_{cls}+\lambda_{r}\mathcal{L}_{loc}$ (4)
3.5 Overall Objective Function
Most existing methods rely on complex combinations of loss functions [46, 47, 23, 55, 39, 50, 33, 58, 34, 40], such as a combination of several or all of the focal loss [38], the scene-class affinity loss [4], the dice loss, the lovasz-softmax loss [2], the depth loss [21], etc., to achieve the expected performance. In contrast, our approach, utilizing the proposed auxiliary supervision for 3D object detection, requires only a simple cross-entropy loss $\mathcal{L}_{ce}$ for occupancy prediction. The total loss for our framework can be defined as:
$\mathcal{L}_{total}=\mathcal{L}_{ce}+\lambda\mathcal{L}_{det}$ (5)
4 Experiments
Method mIoU Modality Input Size 2D Backbone
$\blacksquare$ others
$\blacksquare$ barrier
$\blacksquare$ bicycle
$\blacksquare$ bus
$\blacksquare$ car
$\blacksquare$ const. veh.
$\blacksquare$ motorcycle
$\blacksquare$ pedestrian
$\blacksquare$ traffic cone
$\blacksquare$ trailer
$\blacksquare$ truck
$\blacksquare$ drive. suf.
$\blacksquare$ other flat
$\blacksquare$ sidewalk
$\blacksquare$ terrain
$\blacksquare$ manmade
$\blacksquare$ vegetation
CTF-Occ [41] 28.53 C 928 $\times$ 1600 R101-DCN 8.1 39.3 20.6 38.3 42.2 16.9 24.5 22.7 21.1 23.0 31.1 53.3 33.8 38.0 33.2 20.8 18.0
BEVDetOcc(2f) [12] 42.02 C 512 $\times$ 1408 Swin-B 12.2 49.6 25.1 52.0 54.5 27.9 28.0 28.9 27.2 36.4 42.2 82.3 43.3 54.6 57.9 48.6 43.6
PanoOcc(4f) [47] 42.13 C 864 $\times$ 1600 R101-DCN 11.7 50.5 29.6 49.4 55.5 23.3 33.3 30.6 31.0 34.4 42.6 83.3 44.2 54.4 56.0 45.9 40.4
FB-OCC(16f) [23] 48.90 C 960 $\times$ 1760 VoVNet-99 14.3 57.0 38.3 57.7 62.1 34.4 39.4 38.8 39.4 42.9 50.0 86.0 50.2 60.1 62.5 52.4 45.7
FlashOcc(2f)* [54] 43.52 C 512 $\times$ 1408 Swin-B 13.4 51.1 27.7 51.6 56.2 27.3 30.0 29.9 29.8 37.8 43.5 83.8 46.6 56.2 59.6 50.8 44.7
OSP [39] 39.41 C 900 $\times$ 1600 R101-DCN 11.2 47.3 27.1 47.6 53.7 23.2 29.4 29.7 28.4 32.4 39.9 79.4 41.4 50.3 53.2 40.5 35.4
DHD(2f) [50] 45.53 C 512 $\times$ 1408 Swin-B 14.1 53.1 32.4 52.4 57.4 30.8 35.2 33.0 33.4 37.9 45.3 84.6 48.0 57.4 60.3 52.3 46.2
RadOcc [57] 49.38 C+L 512 $\times$ 1408 Swin-B 10.9 58.2 25.0 57.9 62.9 34.0 33.5 50.1 32.1 48.9 52.1 82.9 42.7 55.3 58.3 68.6 66.0
OccFusion [33] 46.67 C+L+R 900 $\times$ 1600 R101-DCN 12.4 50.3 31.5 57.6 58.8 34.0 41.0 47.2 29.7 42.0 48.0 78.4 35.7 47.3 52.7 63.5 63.3
OccFusion [58] 48.74 C+L 900 $\times$ 1600 R101 12.4 51.8 33.0 54.6 57.7 34.0 43.0 48.4 35.5 41.2 48.6 83.0 44.7 57.1 60.0 62.5 61.3
HyDRa [49] 44.4 C+R 256 $\times$ 704 R50 - - - 52.3 56.3 - 35.9 35.1 - - 44.1 - - - - - -
EFFOcc [40] 52.62 C+L 512 $\times$ 1408 Swin-B 14.7 58.6 34.7 60.5 65.5 36.0 40.4 51.5 41.0 48.5 54.6 84.2 46.4 57.9 60.8 70.9 68.4
DAOcc 53.82 C+L 256 $\times$ 704 R50 12.4 59.6 38.4 61.9 67.1 35.3 48.2 59.1 43.5 50.9 56.3 83.0 44.7 56.7 59.9 70.0 68.1
Table 1: 3D Semantic occupancy prediction performance on Occ3D-nuScenes validation set. (xf) means use x frames for temporal fusion. * denotes the usage of exponantial moving average (EMA). The C, L, and R denote camera, LiDAR, and radar, respectively.
Method RayIoU Modality Input Size 2D Backbone RayIoU1m RayIoU2m RayIoU4m mIoU
BEVFormer(4f) [22] 32.4 C 900 $\times$ 1600 R101 26.1 32.9 38.0 39.2
RenderOcc [35] 19.5 C 512 $\times$ 1408 Swin-B 13.4 19.6 25.5 24.4
SimpleOcc [8] 22.5 C 336 $\times$ 672 R101 17.0 22.7 27.9 31.8
BEVDetOcc(8f) [12] 32.6 C 384 $\times$ 704 R50 26.6 33.1 38.2 39.3
FB-Occ(16f) [23] 33.5 C 256 $\times$ 704 R50 26.7 34.1 39.7 39.1
SparseOcc(16f) [29] 36.1 C 256 $\times$ 704 R50 30.2 36.8 41.2 30.9
Panoptic-FlashOcc [55] 35.2 C 256 $\times$ 704 R50 29.4 36.0 40.1 29.4
Panoptic-FlashOcc(8f) [55] 38.5 C 256 $\times$ 704 R50 32.8 39.3 43.4 31.6
DAOcc 48.2 C+L 256 $\times$ 704 R50 44.4 48.7 51.6 47.88
Table 2: 3D Semantic occupancy prediction performance on Occ3D-nuScenes validation set without camera mask. (xf) means use x frames for temporal fusion. The C and L denote camera, LiDAR, respectively.
Method IoU mIoU Modality Input Size 2D Backbone
$\blacksquare$ barrier
$\blacksquare$ bicycle
$\blacksquare$ bus
$\blacksquare$ car
$\blacksquare$ const. veh.
$\blacksquare$ motorcycle
$\blacksquare$ pedestrian
$\blacksquare$ traffic cone
$\blacksquare$ trailer
$\blacksquare$ truck
$\blacksquare$ drive. suf.
$\blacksquare$ other flat
$\blacksquare$ sidewalk
$\blacksquare$ terrain
$\blacksquare$ manmade
$\blacksquare$ vegetation
MonoScene [4] 24.0 7.3 C 900$\times$ 1600 R101-DCN 4.0 0.4 8.0 8.0 2.9 0.3 1.2 0.7 4.0 4.4 27.7 5.2 15.1 11.3 9.0 14.9
BEVFormer [22] 30.5 16.7 C 900$\times$ 1600 R101-DCN 14.2 6.5 23.4 28.2 8.6 10.7 6.4 4.0 11.2 17.7 37.2 18.0 22.8 22.1 13.8 22.2
SurroundOcc [48] 31.4 20.3 C 900$\times$ 1600 R101-DCN 20.5 11.6 28.1 30.8 10.7 15.1 14.0 12.0 14.3 22.2 37.2 23.7 24.4 22.7 14.8 21.8
OccFormer [59] 29.9 20.1 C 896$\times$ 1600 R101 21.1 11.3 28.2 30.3 10.6 15.7 14.4 11.2 14.0 22.6 37.3 22.4 24.9 23.5 15.2 21.1
C-CONet [46] 26.1 18.4 C 896$\times$ 1600 R101 18.6 10.0 26.4 27.4 8.6 15.7 13.3 9.7 10.9 20.2 33.0 20.7 21.4 21.8 14.7 21.3
FB-Occ [23] 31.5 19.6 C 896$\times$ 1600 R101 20.6 11.3 26.9 29.8 10.4 13.6 13.7 11.4 11.5 20.6 38.2 21.5 24.6 22.7 14.8 21.6
RenderOcc [35] 29.2 19.0 C 896$\times$ 1600 R101 19.7 11.2 28.1 28.2 9.8 14.7 11.8 11.9 13.1 20.1 33.2 21.3 22.6 22.3 15.3 20.9
LMSCNet [37] 36.6 14.9 L - - 13.1 4.5 14.7 22.1 12.6 4.2 7.2 7.1 12.2 11.5 26.3 14.3 21.1 15.2 18.5 34.2
L-CONet [46] 39.4 17.7 L - - 19.2 4.0 15.1 26.9 6.2 3.8 6.8 6.0 14.1 13.1 39.7 19.1 24.0 23.9 25.1 35.7
M-CONet [46] 39.2 24.7 C+L 896$\times$ 1600 R101 24.8 13.0 31.6 34.8 14.6 18.0 20.0 14.7 20.0 26.6 39.2 22.8 26.1 26.0 26.0 37.1
Co-Occ [34] 41.1 27.1 C+L 896$\times$ 1600 R101 28.1 16.1 34.0 37.2 17.0 21.6 20.8 15.9 21.9 28.7 42.3 25.4 29.1 28.6 28.2 38.0
OccFusion [33] 44.7 27.3 C+L+R 900$\times$ 1600 R101-DCN 27.1 19.6 33.7 36.2 21.7 24.8 25.3 16.3 21.8 30.0 39.5 19.9 24.9 26.5 28.9 40.4
DAOcc 45.0 30.5 C+L 256$\times$ 704 R50 30.8 19.5 34.0 38.8 25.0 27.7 29.9 22.5 23.2 31.6 41.0 25.9 29.4 29.9 35.2 43.5
Table 3: 3D Semantic occupancy prediction performance on SurroundOcc validation set. The C, L, and R denote camera, LiDAR, and radar, respectively.
4.1 Datasets and Evaluation Metrics
Occ3D-nuScenes [41] is a large-scale benchmark for 3D occupancy prediction, containing 700 scenes for training and 150 for validation, each lasting 20 seconds with annotations at 2 Hz. The dataset covers a perception range of [-40m, 40m] in the $X$ and $Y$ directions and [-1m, 5.4m] along the $z$-axis, which is discretized into voxels of [0.4m, 0.4m, 0.4m] in size. Each occupied voxel is assigned one of 17 semantic labels, including 16 common classes and a general object class ‘others’. The data includes LiDAR point clouds and RGB images from six surround cameras, enabling dense voxel-wise annotation for 3D scene understanding. In addition, the dataset also provides visibility masks for LiDAR and camera modes, and it can be used for training.
SurroundOcc [48] is an automatically generated occupancy prediction dataset without human annotation. It is built upon nuScenes [3] and generates dense occupancy ground truth utilizing existing 3D detection and 3D semantic segmentation labels. The dataset employs the training set to train the model and utilizes the validation set for evaluation purposes. The occupancy prediction range is set to [-50m, 50m] for the $X$ and $Y$ axes, and [-5m, 3m] for the $z$-axis. The final output occupancy grid has a shape of 200×200×16, with a voxel size of [0.5m, 0.5m, 0.5m].
Evaluation metrics: Following the previous methods, when training Occ3D-nuScenes with the camera visible mask, we use the mean Intersection over Union (mIoU) for fair comparison with previous methods. When the camera visible mask is not used during training, we report an additional semantic segmentation metric, RayIoU [29], which was recently proposed. For SurroundOcc, we report semantic segmentation performance using IoU and mIoU, where IoU is the intersection over union of occupied voxels. This IoU is used as an evaluation metric for the scene completion (SC) task by ignoring the semantic categories of occupied voxels.
4.2 Implementation Details
We implement our method using PyTorch [36] and train all models on 8 NVIDIA RTX 4090 GPUs. We utilize ResNet-50 [11], which is pre-trained on nuImages [3], as the image backbone and crop the input image size to 256×704. In the LiDAR branch, we voxelize 10 LiDAR scans and start training from scratch. We employ the AdamW [32] optimizer with a cosine annealing learning rate scheduler, including a warmup phase, and set an initial learning rate of 2e-4. We apply data augmentation to both the input image and the ground truth 3D bounding boxes, and adopt CBGS [61].
Figure 3: Qualitative visualizations on Occ3D-nuScenes validation set without camera mask.
4.3 Comparison with State-of-the-Art Methods
Comparison on Occ3D-nuScenes with camera mask. As presented in Table 1, we compare our DAOcc with both single-modal and multi-modal methods on Occ3D-nuScenes, all results are either implemented directly by their authors or reported based on their official code implementations. All methods use the camera visible mask during training, except for CTF-Occ [41]. Our DAOcc establishes a new state-of-the-art performance, achieving an outstanding 53.82 mIoU, while utilizing a 2D backbone ResNet50 and 256×704 input image resolution. Compared with the current state-of-the-art multi-modal method EFFOcc [40], our DAOcc outperforms EFFOcc by 1.2 mIoU. Moreover, while EFFOcc employs Swin-B and an input image resolution of 512×1408, DAOcc utilizes only ResNet50 and an input image resolution of 256×704. In addition, EFFOcc requires the LiDAR branch to perform 3D object detection pre-training, which results in an improvement of approximately 3 mIoU, in order to achieve optimal results, whereas DAOcc does not require 3D detection pretraining.
Comparison on Occ3D-nuScenes without camera mask. To the best of our knowledge, no existing multi-modal occupancy prediction work has used RayIoU metrics to report model performance, so we mainly compare with image-based methods. As shown in Table 2, thanks to the multi-modal input and the effectiveness of our proposed method, our DAOcc is significantly ahead of image-based methods, although many of these image-based methods use multi-frame temporal fusion. It should be noted that some image-based methods based on monocular depth estimation, such as Panoptic-FlashOcc [55], only require images as input during the inference phase, but require additional point cloud data input to provide depth information supervision during the training phase, so these methods can also be regarded as multi-modal during the training phase.
Comparison on SurroundOcc. Table 3 presents a quantitative comparison on SurroundOcc validation set, showcasing the performance of our DAOcc against other methods. By leveraging both LiDAR and camera inputs, DAOcc achieves a remarkable 3.2 mIoU increase over OccFusion [33], which not only employs camera and LiDAR but also incorporates radars. Notably, OccFusion utilizes R101-DCN with a high-resolution input of 900×1600, whereas our method employs a more lightweight Resnet50 backbone and a lower resolution of 256×704. This underscores the effectiveness and efficiency of our proposed method.
4.4 Ablation Studies
Ablation studies are conducted on Occ3D-nuScenes using the camera mask. Unless otherwise specified, all experiments presented in this section are trained for 15 epochs only.
BEV View Range Voxel Resolution (BEV) mIoU
(a) $\left[-41.4,41.4\right]$ 0.075 $\times$ 0.075 50.76
(b) $\left[-51.2,51.2\right]$ 0.1 $\times$ 0.1 49.98
(c) $\left[-54.0,54.0\right]$ 0.075 $\times$ 0.075 51.26
Table 4: Performance comparison between different BEV processing ranges and voxel sizes.
We propose the BVRE strategy to enrich the spatial contextual semantic information by broadening the perceptual range from the BEV perspective. As shown in Table 4, comparing (a) and (c), increasing the BEV view range from [-41.4, 41.4] to [-54.0, 54.0] can bring a performance improvement of 0.5 mIoU. In addition, by comparing (a) and (b), we can find that the size of the voxel has a great impact on the model’s performance. Setting a smaller voxel size will bring obvious performance improvements but will also impose a greater computational burden.
Detection Auxiliary mIoU bicycle motorcycle traffic cone
51.26 20.60 38.89 36.65
52.82 36.35 45.48 41.98
Table 5: Ablation study on the auxiliary detection head.
To fully leverage the inherent geometric structures within point cloud features, we incorporate 3D object detection as auxiliary supervision. As shown in Table 5, by adding auxiliary supervision for 3D object detection, we can achieve an improvement of 1.56 mIoU, demonstrating the effectiveness of our method. Given that the gain in object detection for foreground objects will be greater, we present the three categories with the highest gain in foreground objects, which are the bicycle (+15.75 mIoU), the motorcycle (+6.59 mIoU) and the traffic cone (+5.33 mIoU). This further proves the rationality of the performance improvement of our method.
Detection Auxiliary EP15 EP24 CBGS (EP6) mIoU
52.82
53.39
53.82
Table 6: Ablation for obtaining the final performance.
Table 6 provides ablation experiments to obtain the final performance. First, by increasing the training time from 15 epochs to 24 epochs, an improvement of 0.57 mIoU can be achieved. Then, by using CBGS [61], another improvement of 0.43 mIoU can be achieved. When using CBGS, since resampling increases the number of samples in each epoch, for the sake of a fair comparison, we limited the training to 6 epochs to approximate the total training time of 24 epochs without CBGS.
4.5 Visualization
Figure 3 presents the visualizations of DAOcc on Occ3D-nuScenes validation set without using camera mask. The results show that our DAOcc made a relatively complete prediction of the scene, accurately reproducing it in fine detail.
5 Conclusion
In this paper, we propose a novel multi-modal occupancy prediction framework, aiming to achieve superior performance while using a deployment-friendly image feature extraction network and input image resolution. Extensive experiments on Occ3D-nuScenes and SurroundOcc demonstrate the superiority of our approach over existing methods. We believe that DAOcc can be put into practice and inspire future research.

### v2

@@
 Multi-sensor fusion significantly enhances the accuracy and robustness of 3D semantic occupancy prediction, which is crucial for autonomous driving and robotics.
-However, existing approaches depend on large image resolutions and complex networks to achieve top performance, hindering their application in practical scenarios.
+However, most existing approaches depend on large image resolutions and complex networks to achieve top performance, hindering their application in practical scenarios.
 Additionally, most multi-sensor fusion approaches focus on improving fusion features while overlooking the exploration of supervision strategies for these features.
-To this end, we propose DAOcc, a novel multi-sensor fusion occupancy network that leverages 3D object detection supervision to assist in achieving superior performance, while using a deployment-friendly image feature extraction network and practical input image resolution.
+To this end, we propose DAOcc, a novel multi-modal occupancy prediction framework that leverages 3D object detection supervision to assist in achieving superior performance, while using a deployment-friendly image feature extraction network and practical input image resolution.
 Furthermore, we introduce a BEV View Range Extension strategy to mitigate the adverse effects of reduced image resolution.
-As a result, our approach achieves new state-of-the-art results on the Occ3D-nuScenes and SurroundOcc datasets, using ResNet50 and a 256x704 input image resolution.
+Experimental results show that DAOcc achieves new state-of-the-art performance on the Occ3D-nuScenes and SurroundOcc benchmarks, and surpasses other methods by a significant margin while using only ResNet50 and 256×704 input image resolution.
 Code will be made available at https://github.com/AlphaPlusTT/DAOcc.
 1 Introduction
-Figure 1: Unlike existing multi-modal methods [46, 33, 58, 40, 34], which rely on complex combinations of loss functions, including several or all of the focal loss $\mathcal{L}_{focal}$ [38], the scene-class affinity loss $\mathcal{L}_{scal}^{geo}$ and $\mathcal{L}_{scal}^{sem}$ [4], the lovasz-softmax loss $\mathcal{L}_{ls}$ [2], and the depth loss $\mathcal{L}_{d}$ [21], to achieve optimal performance, our approach only requires a single cross-entropy loss $\mathcal{L}_{ce}$ for the occupancy prediction task.
+Figure 1: Comparison of loss functions.
+Unlike existing multi-modal methods [49, 37, 61, 43, 38], which rely on complex combinations of loss functions, including several or all of the focal loss $\mathcal{L}_{focal}$ [41], the scene-class affinity loss $\mathcal{L}_{scal}^{geo}$ and $\mathcal{L}_{scal}^{sem}$ [4], the lovasz-softmax loss $\mathcal{L}_{ls}$ [2], and the depth loss $\mathcal{L}_{d}$ [22], to achieve optimal performance, our approach only requires a single cross-entropy loss $\mathcal{L}_{ce}$ for the occupancy prediction task.
 3D semantic occupancy prediction (occ) is a critical task in autonomous driving [54, 4, 50] and robotic systems [45, 47, 15, 48], where accurately understanding the environment is essential for safe and efficient navigation.
@@
 However, this approach significantly limits the deployment of top-performing occ models on edge devices due to the high computational demands.
-Compared to image, point cloud is much sparser.
-For example, in the training set of the nuScenes dataset, the maximum number of points in a single frame point cloud is only 34,880, which is equivalent to just 2.4% of the number of pixels in a 900×1600 resolution image.
+In comparison to images, point cloud is much sparser.
+For example, in the training set of the nuScenes [3] dataset, the maximum number of points in a single frame point cloud is only 34,880, which is equivalent to just 2.4% of the number of pixels in a 900×1600 resolution image.
 Therefore, how to effectively leverage point cloud data within a multi-modal occ framework remains to be further explored.
 Moreover, we observe that most works [60, 37, 61, 52, 43, 49] on multi-modal occ primarily focuses on obtaining more effective fusion features, with insufficient exploration into the forms of supervision for these fused features.
-Although CO-Occ [34] introduces regularization based on implicit volumetric rendering to supervise fused features, it only utilizes distance ground truth from the original point cloud data, failing to fully exploit the geometric and structural information inherent in point cloud features.
-In contrast, point cloud-based 3D object detectors [56, 6, 31] effectively leverage this information, achieving significantly better performance in 3D object detection tasks compared to image-based detectors [26, 16, 27, 28].
+Although CO-Occ [38] introduces regularization based on implicit volumetric rendering to supervise fused features, it only utilizes distance ground truth from the original point cloud data, failing to fully exploit the geometry and structure information inherent in point cloud features.
+In contrast, LiDAR-based 3D detectors [59, 7, 34] effectively leverage this information, achieving significantly better performance in 3D object detection tasks compared to image-based 3D detectors [28, 17, 29, 30].
 This observation suggests a new research direction: how to better exploit the unique strengths of point cloud data in multi-modal occ tasks.
 Figure 2: Overview of our proposed DAOcc.
-We first introduce the BVRE strategy to enrich the spatial contextual semantic information by broadening the perceptual range from the BEV perspective.
+We first introduce the BVRE strategy to enrich the spatial contextual information by broadening the perceptual range from the BEV perspective.
 Then feature extraction is performed through the multi-modal occupancy prediction network illustrated in the above figure.
-Furthermore, to fully leverage the inherent geometric structures within point cloud features, we incorporate 3D object detection as auxiliary supervision.
-Based on these observations, we propose DAOcc, a novel multi-modal occ framework that leverages 3D object detection to assist in achieving superior performance while using a deployment-friendly image feature extraction network and practical input image resolution.
-In constructing the baseline network structure of DAOcc, we adopt the most direct and straightforward approach: first, we extract features from images and point clouds, respectively.
-Given that depth estimation from monocular images is an ill-posed problem [13, 20] and that deformable attention modules are overly complex [10], we employ a simple method, similar to Harley et al.
-[9], to transform image features from 2D space to 3D volumetric space.
-Specifically, we project a set of predefined 3D points onto the 2D image feature plane and use bilinear interpolation to sample the corresponding 2D image features for these 3D points.
-Next, we adopt the same simplest fusion strategy as BEVFusion.
-[30], concatenating image and point cloud features, then performing a 2D convolution to fuse them, leading to a unified BEV features.
-Finally, we apply a fully convolutional BEV encoder with a residual structure to further fuse the unified BEV features, and subsequently employ the Channel-to-Height operation [54] to reshape the height of the unified BEV features.
-To fully exploit the geometric and structural information inherent in point cloud features, we augment the baseline model’s unified BEV features with 3D object detection supervision, thereby enhancing the discriminability of the unified BEV features.
-This renders the unified BEV features more sensitive to object boundaries and capable of perceiving relationships between internal object structures.
+Furthermore, to fully leverage the inherent geometry and structure information within point cloud features, we incorporate 3D object detection as auxiliary supervision.
+Based on these observations, we propose DAOcc, a novel multi-modal occ framework that leverages 3D object detection to assist in achieving superior performance while using a deployment-friendly image encoder and practical input image resolution.
+In constructing the baseline network of DAOcc, we adopt the simplest and most direct approach.
+First, we use a 2D image encoder and 3D sparse convolutions [55] to extract features from images and point clouds, respectively.
+Given that depth estimation from monocular images is an ill-posed problem [14, 21] and that deformable attention modules are overly complex [11], we employ a simple method similar to Harley et al.
+[10] to transform image features from 2D space to 3D volumetric space.
+Specifically, we project a set of predefined 3D points onto the 2D image feature plane and use bilinear interpolation to sample the corresponding 2D image features for these points.
+Next, we adopt a straightforward fusion strategy by concatenating the image and point cloud features and then applying 2D convolution to obtain a unified BEV (Bird’s Eye View) feature.
+Finally, we apply a fully convolutional BEV encoder with a residual structure to further fuse the unified BEV features, and subsequently employ the Channel-to-Height [57] operation to transform the channel dimension into the height dimension.
+To fully exploit the geometry and structure information inherent in point cloud features, we augment the baseline model’s unified BEV features with 3D object detection supervision, thereby enhancing the discriminability of the unified BEV features.
+This renders the unified BEV features more sensitive to object boundaries and enables our method to avoid the need for complex loss function combinations, as depicted in Figure 1.
 Additionally, given the sparsity of point clouds, we extend the processing range of the point clouds and employ sparse convolutions [55] to mitigate the computational overhead introduced by this extension.
@@
 It is worth noting that the 3D object detection supervision is only used as an auxiliary branch during training and can be removed during inference for the occ task.
-As a result, the proposed DAOcc establishes the new state-of-the-art performance on the Occ3D-nuScenes and SurroundOcc benchmarks, while using ResNet50 and a 256×704 input image resolution.
+Extensive experiments are conducted to validate the efficiency and effectiveness of the proposed DAOcc.
+The results show that DAOcc establishes the new state-of-the-art performance on the Occ3D-nuScenes [44] and SurroundOcc [51] benchmarks, while using only ResNet50 [12] and a 256x704 input image resolution.
 Specifically, on the Occ3D-nuScenes validation set, DAOcc achieves 53.82 mIoU when using the camera mask during training and 48.2 RayIoU without using the camera mask.
@@
 In conclusion, our contributions are summarized as follows:
-  • We design a simple yet efficient multi-modal 3D semantic occupancy prediction baseline, eliminating the need for complex deformable attention modules as well as image depth estimation during feature fusion.
-  • We propose DAOcc, a novel multi-modal occupancy prediction framework that leverages 3D object detection to assist in achieving superior performance while using a deployment-friendly image feature extraction network and input image resolution.
+  • We design a simple yet efficient multi-modal 3D semantic occupancy prediction baseline, eliminating the need for complex deformable attention [23] modules as well as image depth estimation during feature fusion.
+  • We propose DAOcc, a novel multi-modal occupancy prediction framework that leverages 3D object detection to assist in achieving superior performance while using a deployment-friendly image encoder and practical input image resolution.
   • We introduce a BEV View Range Extension strategy, which provides a larger BEV field of view, offering more contextual information and mitigating the adverse effects of reduced image resolution.
-  • We establish the new state-of-the-art performance on Occ3D-nuScenes dataset and SurroundOcc dataset, while using ResNet50 and a 256×704 input image resolution.
+  • We establish the new state-of-the-art performance on Occ3D-nuScenes dataset and SurroundOcc dataset, while using only ResNet50 and a 256×704 input image resolution.
 2 Related Work
@@
 Camera-based 3D occupancy prediction has gained significant attention due to its cost-effectiveness.
-MonoScene [4] is the first work to infer the dense occupancy and semantics of outdoor and indoor scenes from a single RGB image.
-Since full perception of the surrounding environment is crucial for making accurate decisions in autonomous driving and robotics, most of the recent work is based on surround view image input.
-TPVFormer [15] proposes an efficient tri-perspective view (TPV) representation that combines BEV with two additional perpendicular planes to provide a 3D perception result with multi-view image inputs.
-Given the insufficient fine-grained semantic information of the TPV representation, OccFormer [59] leverages dense 3D features and proposes a dual-path transformer-based occupancy network.
-However, both TPVFormer and OccFormer use sparse LiDAR points as supervision, resulting in sparse occupancy predictions.
+Early works [4, 16, 62] lack dense surround occupancy supervision, resulting in sparse occupancy predictions.
 To obtain dense occupancy prediction, OpenOccupancy [49], SurroundOcc [51] and Occ3D [44] developed methods for generating dense occupancy labels and established benchmarks on their respective proposed datasets.
@@
 PanoOcc [50] adopts voxel query and a coarse-to-fine approach to learn a unified occupancy representation.
-FlashOcc [54] uses BEV features with 2D convolution to achieve efficient feature extraction and proposes the channel-to-height transformation to lift the output of BEV to 3D space.
+FlashOcc [57] uses BEV features with 2D convolution to achieve efficient feature extraction and proposes the Channel-to-Heigh transformation to lift the output of BEV to 3D space.
 SparseOcc [31] leverages a sparse voxel decoder to reconstruct the sparse geometry of a scene and uses sparse queries to predict masks and labels.
-OSP [39] treats the scene as a collection of points and uses flexible sampling to allow the model to pay special attention to certain areas.
 However, all camera-based methods inevitably need to cope with the challenges of adverse lighting and weather conditions, and many works inevitably lack robustness due to the ill-posed monocular depth estimation problem [14, 21].
@@
 In concurrent work, OccFusion [61] projects the preprocessed denser and more uniform point cloud onto the image plane to establish a mapping relationship, and performs deformable attention [67] to fuse the corresponding features.
-Although OccFusion [58] avoids depth estimation, using deformable attention incurs a greater computational burden.
+Although OccFusion [61] avoids depth estimation, using deformable attention incurs a greater computational burden [11].
 Hydra [52] extends FB-BEV [25] to the camera-radar fusion domain and improves the accuracy of depth estimation with the help of radar features.
@@
 However, this regularization only utilizes the distance ground truth of the point cloud and does not leverage its intrinsic geometry information.
-EFFOcc [40] proposes an efficient and lightweight multi-modal 3D occupancy network, but it significantly relies on the 3D detection pretraining of the LiDAR branch for optimal results.
+EFFOcc [43] proposes an efficient and lightweight multi-modal 3D occupancy network, but it significantly relies on the 3D object detection pre-training of the LiDAR branch for optimal results.
 This dependence, however, limits the flexibility of the network structure design.
-Specifically, to maximize the benefits of the 3D detection pretraining, EFfOcc has to adjust its network architecture to align more closely with well-established LiDAR detection networks.
-In addition, the inherently ill-posed nature of the monocular depth estimation problem inevitably renders the image branch of EFFOcc less robust [13].
-In this work, we introduce a simple yet efficient multi-modal occupancy prediction network, eliminating the need for complex deformable attention [63, 22] as well as image depth estimation during feature fusion, and does not require 3D object detection pretraining.
-Furthermore, to fully exploit the inherent geometry information in point cloud features, we add supervision of 3D object detection as an auxiliary branch on the fused features.
+Specifically, to maximize the benefits of the 3D detection pre-training, EFFOcc has to adjust its network architecture to align more closely with well-established LiDAR detection networks.
+In this work, we introduce a simple yet efficient multi-modal occupancy prediction network, eliminating the need for complex deformable attention [23] as well as image depth estimation during feature fusion, and does not require 3D object detection pre-training.
+Furthermore, to fully exploit the inherent geometry information in point cloud features, we add supervision of 3D object detection as an auxiliary supervision on the fused features.
 Multi-Modal 3D Object Detection: Recent multi-modal 3D object detection methods [33, 19, 20, 6, 46] mainly focus on learning effective BEV feature representations.
 TransFusion [1] proposes a two-stage transformer-decoder based detection head and applies cross attention to obtain image features for each object query.
-BEVFusion [30] proposes an efficient and generic multi-task multi-sensor fusion framework which unifies multi-modal features in the shared bird’s-eye view (BEV) representation space, and introduces a specialized kernel to speed up the BEV pooling operation.
-In concurrent work, to improve the robustness of the LiDAR-camera fusion framework to sensor failures, such as missing LiDAR sensor input, BEVFusion decomposes the LiDAR-camera fusion into two streams that can independently output perception results, and performs feature fusion after the two streams.
+BEVFusion [33] proposes an efficient and generic multi-task multi-sensor fusion framework which unifies multi-modal features in the shared BEV representation space, and introduces a specialized kernel to speed up the BEV pooling operation.
+In concurrent work, BEVFusion [26] decomposes the LiDAR-camera fusion into two streams that can independently output perception results, and performs feature fusion after the two streams.
 DAL [14] follows the concept of ‘Detecting As Labeling’ and decouples the fused features during classification and regression.
 Specifically, it employs the fused features for classification, whereas it utilizes the point cloud features exclusively for regression.
-In this work, we incorporate the simple yet effective feature fusion approach of BEVFusion and introduce 3D object detection as an auxiliary branch during training.
+In this work, we incorporate the simple yet effective feature fusion approach of BEVFusion [33] and introduce 3D object detection as an auxiliary branch during training.
 3 Proposed Method
 3.1 Overall Framework
-Our goal is to fully exploit the advantages of point cloud features for multi-modal occupancy prediction.
-Previous multi-modal studies have not fully exploited this and can only achieve superior performance through more complex image feature extraction networks and larger input image resolutions.
+Our purpose is to fully utilize the inherent geometry and structure information in point cloud features within a multi-modal occupancy prediction framework.
+Previous multi-modal methods have not fully exploited this and can only achieve superior performance through a more complex image encoder and a larger input image resolution.
 The overall framework of our proposed DAOcc is illustrated in Figure 2.
 DAOcc takes surround images and their corresponding time-synchronized point cloud as input, and obtains the features of the image and point cloud through the Camera Encoder and LiDAR Encoder, respectively.
-The 2D image features are transformed into the 3D voxel space by projection and sampling.
-Subsequently, the image and point cloud features in the 3D space are compressed along the height dimension to generate the corresponding BEV features.
+The 2D image features are transformed into the 3D voxel space by projection and interpolation.
+Subsequently, the image and point cloud features in the 3D space are compressed along the height dimension to generate the corresponding BEV(Bird’s Eye View) features.
 A simple 2D convolution is then applied for feature fusion, and the fully convolutional BEV Encoder encodes these fused features to obtain the final BEV representations.
 Finally, the occupancy head uses the Channel-to-Height operation [57] to restore the height of the BEV representations, resulting in final 3D voxel space representations that can be utilized for occupancy prediction.
-These modules collectively form the basic network architecture of DAOcc, which will be elaborated in Section 3.2.
-On the foundation of the basic network, we introduce the BVRE strategy (see Section 3.3) to compensate for the information loss that arises from the reduction in image resolution.
-This strategy aims to enrich the spatial contextual semantic information by broadening the perceptual range from the Bird’s Eye View (BEV) perspective.
-Furthermore, to fully leverage the inherent geometric structures within point cloud features, we incorporate 3D object detection as auxiliary supervision (see Section 3.4).
-This auxiliary supervision not only enhances the discriminability of the fused features but also leads to a very concise overall training loss (see Section 3.5) for our proposed framework.
+These modules collectively form the basic network architecture of DAOcc, which will be elaborated in Sec.
+3.2.
+On the foundation of the basic network, we introduce the BVRE strategy (Sec.
+3.3) to compensate for the information loss that arises from the reduction in image resolution.
+This strategy aims to enrich the spatial contextual semantic information by broadening the perceptual range from the BEV perspective.
+Furthermore, to fully leverage the inherent geometry and structure information within point cloud features, we incorporate 3D object detection as auxiliary supervision (Sec.
+3.4).
+This auxiliary supervision not only enhances the discriminability of the fused features but also leads to a very concise overall training loss (Sec.
+3.5) for our proposed framework.
 3.2 Basic Network
 LiDAR Encoder.
-The method of embedding raw LiDAR points into 3D voxelized features is consistent with BEVFusion [30].
+The method of embedding raw LiDAR points into 3D voxelized features is consistent with SECOND [55].
 We first voxelize the point cloud, retaining a maximum of 10 points per voxel, resulting in a 3D voxel grid of size $D\times H\times W$.The feature representation for each voxel is obtained by averaging the features of all points within it.
@@
 Camera Encoder.
-For image feature extraction, taking surround images as input, we first use ResNet50 [11] as the backbone to extract multi-scale features, denoted as $F_{ms}=\{F_{ms\times\frac{1}{8}},F_{\frac{1}{16}},F_{\frac{1}{32}}\}$, where $F_{\frac{1}{x}}$ represent features extracted after $x\times$ downsampling.
+For image feature extraction, taking surround images as input, we first use ResNet50 [12] as the backbone to extract multi-scale features, denoted as $F_{ms}=\{F_{\frac{1}{8}},F_{\frac{1}{16}},F_{\frac{1}{32}}\}$, where $F_{\frac{1}{x}}$ represent features extracted after $x\times$ downsampling.
 Then, we employ the Feature Pyramid Network (FPN) [27] as the neck to aggregate these multi-scale features.
@@
 For voxels located in the overlapping viewing regions of the surround cameras, we average the image features from the two corresponding cameras to obtain the final feature for each voxel.
-The output camera voxel features can be denoted as $F_{c}\in\mathbb{R}^{C\times Z\times\frac{H}{8}\times\frac{W}{8}}$.
+The output camera’s voxel features can be denoted as $F_{c}\in\mathbb{R}^{C\times Z\times\frac{H}{8}\times\frac{W}{8}}$.
 BEV Encoder.
@@
 The output refined BEV features can be denoted as $F_{r}\in\mathbb{R}^{C_{r}\times\frac{H}{8}\times\frac{W}{8}}$.
+Figure 3: Comparison of the impact of resolution increase on computational cost.
+For LiDAR Encoder, the processing range of the point cloud along the $z$-axis in all experiments is from -5m to 3m.
+The example in the figure above selects the point cloud with the highest point count in the validation set and aggregates a total of 10 frames.
 3.3 BVRE
-Considering the sparsity of point clouds, the increase in computational cost resulting from expanding the processing range of point clouds is very small compared to images.
+As shown in Figure 3, due to the sparsity of point clouds, the increase in computational cost resulting from expanding the processing range of the point cloud is very small compared to increasing the image resolution.
 Therefore, we extend the point cloud range to provide more 3D spatial context, compensating for the information loss caused by lower image resolution.
 However, arbitrarily setting the $XY$ range of the point cloud or voxel resolution may result in misalignment between the BEV features and the occupancy ground truth.
-Assuming that the voxel resolution of the occupancy grid is $res_{o}$, the expanded point cloud data in BEV covers a rectangular region with $x$-coordinates ranging from $-x$ to $x$ and $y$-coordinates ranging from $-y$ to $y$, and has a voxel resolution of $res_{p}$.
+Assuming that the voxel resolution of the occupancy grid is $res_{o}$, the expanded point cloud data in BEV covers a rectangular region with $X$ ranging from $-x$ to $x$ and $Y$ ranging from $-y$ to $y$, and has a voxel resolution of $res_{p}$.
 Thus, $x$ and $y$ must be integer multiples of $res_{o}$, and $res_{o}$ must also be an integer multiple of $res_{p}$.
-Otherwise, the spatial resolution along the $z$-axis for each feature vector in the BEV representation will not be equal to $res_{o}$.
-To avoid complex manual design, we adopted coordinate transformation and interpolation, as illustrated in the purple part of Figure 2, which can be formulated as follows:
-$F_{occ}={\rm GridSample}(F_{r},{\rm Norm}(T_{o2p}\times P_{o}))$ (1)
-where $P_{o}$ is a set of predefined points in the $XY$ plane of the coordinate system of the occupancy annotation, each point is located at the center of an occupancy grid in the $XY$ plane.
-$T_{o2p}$ is the transformation matrix from occupancy coordinates to Lidar coordinates.
+Otherwise, the spatial resolution represented by each feature vector along the $z$-axis in the BEV representation will not be equal to $res_{o}$.
+To avoid complex manual design, we adopt coordinate transformation and interpolation, as illustrated in the purple part of Figure 2, which can be formulated as follows:
+$F_{occ}={\rm GridSample}(F_{r},{\rm Norm}(T_{o2l}\times P_{o}))$ (1)
+where $P_{o}$ is a set of predefined points in the $XY$ plane of the occupancy annotation coordinate system, each point is located at the center of an occupancy grid in the $XY$ plane.
+$T_{o2l}$ is the transformation matrix from occupancy coordinates to LiDAR coordinates.
 The function of ${\rm Norm}$ is to scale the coordinate values to a range from -1 to 1.
-3.4 Detachable Auxiliary Head
-To fully exploit the inherent geometry information in point clouds, we add a separable auxiliary detection network to the fused features, which will enable the features to benefit more from the geometry information.
+3.4 Detachable Auxiliary Detection Head
+To fully exploit the inherent geometry and structure information in point cloud features, we employ a detachable auxiliary detection head to further supervise the feature fusion process, thereby enhancing the information about object boundaries and scene structures in the fused features.
 Meanwhile, the auxiliary detection task is related to the occupancy prediction task, thus providing multiple regularization effects during optimization.
@@
 The total loss for the auxiliary detection head can be denoted as:
-$\mathcal{L}_{det}=\mathcal{L}_{cls}+\lambda_{r}\mathcal{L}_{loc}$ (4)
+$\mathcal{L}_{det}=\mathcal{L}_{cls}+\lambda_{l}\mathcal{L}_{loc}$ (4)
 3.5 Overall Objective Function
 Most existing methods rely on complex combinations of loss functions [49, 50, 24, 58, 42, 53, 37, 61, 38, 43], such as a combination of several or all of the focal loss [41], the scene-class affinity loss [4], the dice loss, the lovasz-softmax loss [2], the depth loss [22], etc., to achieve the expected performance.
-In contrast, our approach, utilizing the proposed auxiliary supervision for 3D object detection, requires only a simple cross-entropy loss $\mathcal{L}_{ce}$ for occupancy prediction.
+In contrast, our approach utilizes a detachable 3D object detection branch for auxiliary supervision and requires only a simple cross-entropy loss $\mathcal{L}_{ce}$ for occupancy prediction.
 The total loss for our framework can be defined as:
@@
 4 Experiments
-Method mIoU Modality Input Size 2D Backbone
+Method mIoU Modality Image Size 2D Backbone
 $\blacksquare$ others
@@
 $\blacksquare$ vegetation
-CTF-Occ [41] 28.53 C 928 $\times$ 1600 R101-DCN 8.1 39.3 20.6 38.3 42.2 16.9 24.5 22.7 21.1 23.0 31.1 53.3 33.8 38.0 33.2 20.8 18.0
 BEVDetOcc(2f) [13] 42.02 C 512 $\times$ 1408 Swin-B 12.2 49.6 25.1 52.0 54.5 27.9 28.0 28.9 27.2 36.4 42.2 82.3 43.3 54.6 57.9 48.6 43.6
@@
 FB-OCC(16f) [24] 48.90 C 960 $\times$ 1760 VoVNet-99 14.3 57.0 38.3 57.7 62.1 34.4 39.4 38.8 39.4 42.9 50.0 86.0 50.2 60.1 62.5 52.4 45.7
-FlashOcc(2f)* [54] 43.52 C 512 $\times$ 1408 Swin-B 13.4 51.1 27.7 51.6 56.2 27.3 30.0 29.9 29.8 37.8 43.5 83.8 46.6 56.2 59.6 50.8 44.7
-OSP [39] 39.41 C 900 $\times$ 1600 R101-DCN 11.2 47.3 27.1 47.6 53.7 23.2 29.4 29.7 28.4 32.4 39.9 79.4 41.4 50.3 53.2 40.5 35.4
-DHD(2f) [50] 45.53 C 512 $\times$ 1408 Swin-B 14.1 53.1 32.4 52.4 57.4 30.8 35.2 33.0 33.4 37.9 45.3 84.6 48.0 57.4 60.3 52.3 46.2
+FlashOcc(2f) [57] 43.52 C 512 $\times$ 1408 Swin-B 13.4 51.1 27.7 51.6 56.2 27.3 30.0 29.9 29.8 37.8 43.5 83.8 46.6 56.2 59.6 50.8 44.7
+COTR(8f) [36] 46.21 C 512 $\times$ 1408 Swin-B 14.9 53.3 35.2 50.8 57.3 35.4 34.1 33.5 37.1 39.0 45.0 84.5 48.7 57.6 61.1 51.6 46.7
 RadOcc [60] 49.38 C+L 512 $\times$ 1408 Swin-B 10.9 58.2 25.0 57.9 62.9 34.0 33.5 50.1 32.1 48.9 52.1 82.9 42.7 55.3 58.3 68.6 66.0
@@
 EFFOcc [43] 52.62 C+L 512 $\times$ 1408 Swin-B 14.7 58.6 34.7 60.5 65.5 36.0 40.4 51.5 41.0 48.5 54.6 84.2 46.4 57.9 60.8 70.9 68.4
-DAOcc 53.82 C+L 256 $\times$ 704 R50 12.4 59.6 38.4 61.9 67.1 35.3 48.2 59.1 43.5 50.9 56.3 83.0 44.7 56.7 59.9 70.0 68.1
-Table 1: 3D Semantic occupancy prediction performance on Occ3D-nuScenes validation set.
+DAOcc (Ours) 53.82 C+L 256 $\times$ 704 R50 12.4 59.6 38.4 61.9 67.1 35.3 48.2 59.1 43.5 50.9 56.3 83.0 44.7 56.7 59.9 70.0 68.1
+Table 1: 3D occupancy prediction performance on Occ3D-nuScenes validation set.
+The camera visible mask is used during the training phase.
 (xf) means use x frames for temporal fusion.
-* denotes the usage of exponantial moving average (EMA).
 The C, L, and R denote camera, LiDAR, and radar, respectively.
-Method RayIoU Modality Input Size 2D Backbone RayIoU1m RayIoU2m RayIoU4m mIoU
+Image Size and 2D Backbone represent the input image resolution and the image feature extractor, respectively.
+The best results are shown in bold.
+Method RayIoU Modality Image Size 2D Backbone RayIoU1m RayIoU2m RayIoU4m mIoU
 BEVFormer(4f) [23] 32.4 C 900 $\times$ 1600 R101 26.1 32.9 38.0 39.2
@@
 Panoptic-FlashOcc(8f) [58] 38.5 C 256 $\times$ 704 R50 32.8 39.3 43.4 31.6
-DAOcc 48.2 C+L 256 $\times$ 704 R50 44.4 48.7 51.6 47.88
-Table 2: 3D Semantic occupancy prediction performance on Occ3D-nuScenes validation set without camera mask. (xf) means use x frames for temporal fusion. The C and L denote camera, LiDAR, respectively.
-Method IoU mIoU Modality Input Size 2D Backbone
+DAOcc (Ours) 48.2 C+L 256 $\times$ 704 R50 44.4 48.7 51.6 47.9
+Table 2: 3D occupancy prediction performance on Occ3D-nuScenes validation set.
+The camera visible mask is not used during the training phase.
+(xf) means use x frames for temporal fusion.
+The C and L denote camera, LiDAR, respectively.
+The best results are shown in bold.
+Method IoU mIoU Modality Image Size 2D Backbone
 $\blacksquare$ barrier
@@
 $\blacksquare$ vegetation
-MonoScene [4] 24.0 7.3 C 900$\times$ 1600 R101-DCN 4.0 0.4 8.0 8.0 2.9 0.3 1.2 0.7 4.0 4.4 27.7 5.2 15.1 11.3 9.0 14.9
-BEVFormer [22] 30.5 16.7 C 900$\times$ 1600 R101-DCN 14.2 6.5 23.4 28.2 8.6 10.7 6.4 4.0 11.2 17.7 37.2 18.0 22.8 22.1 13.8 22.2
 SurroundOcc [51] 31.4 20.3 C 900$\times$ 1600 R101-DCN 20.5 11.6 28.1 30.8 10.7 15.1 14.0 12.0 14.3 22.2 37.2 23.7 24.4 22.7 14.8 21.8
@@
 OccFusion [37] 44.7 27.3 C+L+R 900$\times$ 1600 R101-DCN 27.1 19.6 33.7 36.2 21.7 24.8 25.3 16.3 21.8 30.0 39.5 19.9 24.9 26.5 28.9 40.4
-DAOcc 45.0 30.5 C+L 256$\times$ 704 R50 30.8 19.5 34.0 38.8 25.0 27.7 29.9 22.5 23.2 31.6 41.0 25.9 29.4 29.9 35.2 43.5
-Table 3: 3D Semantic occupancy prediction performance on SurroundOcc validation set. The C, L, and R denote camera, LiDAR, and radar, respectively.
+DAOcc (Ours) 45.0 30.5 C+L 256$\times$ 704 R50 30.8 19.5 34.0 38.8 25.0 27.7 29.9 22.5 23.2 31.6 41.0 25.9 29.4 29.9 35.2 43.5
+Table 3: 3D occupancy prediction performance on SurroundOcc validation set. The C, L, and R denote camera, LiDAR, and radar, respectively. The best results are shown in bold.
 4.1 Datasets and Evaluation Metrics
-Occ3D-nuScenes [41] is a large-scale benchmark for 3D occupancy prediction, containing 700 scenes for training and 150 for validation, each lasting 20 seconds with annotations at 2 Hz.
-The dataset covers a perception range of [-40m, 40m] in the $X$ and $Y$ directions and [-1m, 5.4m] along the $z$-axis, which is discretized into voxels of [0.4m, 0.4m, 0.4m] in size.
-Each occupied voxel is assigned one of 17 semantic labels, including 16 common classes and a general object class ‘others’.
-The data includes LiDAR point clouds and RGB images from six surround cameras, enabling dense voxel-wise annotation for 3D scene understanding.
-In addition, the dataset also provides visibility masks for LiDAR and camera modes, and it can be used for training.
-SurroundOcc [48] is an automatically generated occupancy prediction dataset without human annotation.
-It is built upon nuScenes [3] and generates dense occupancy ground truth utilizing existing 3D detection and 3D semantic segmentation labels.
-The dataset employs the training set to train the model and utilizes the validation set for evaluation purposes.
-The occupancy prediction range is set to [-50m, 50m] for the $X$ and $Y$ axes, and [-5m, 3m] for the $z$-axis.
-The final output occupancy grid has a shape of 200×200×16, with a voxel size of [0.5m, 0.5m, 0.5m].
-Evaluation metrics: Following the previous methods, when training Occ3D-nuScenes with the camera visible mask, we use the mean Intersection over Union (mIoU) for fair comparison with previous methods.
-When the camera visible mask is not used during training, we report an additional semantic segmentation metric, RayIoU [29], which was recently proposed.
-For SurroundOcc, we report semantic segmentation performance using IoU and mIoU, where IoU is the intersection over union of occupied voxels.
-This IoU is used as an evaluation metric for the scene completion (SC) task by ignoring the semantic categories of occupied voxels.
+Occ3D-nuScenes [44] is a large-scale benchmark dataset for 3D occupancy prediction, which contains 700 scenes for training and 150 scenes for validation.
+Each scene lasts for 20 seconds, with annotations provided at a frequency of 2 Hz.
+The dataset covers a perception range of [-40m, 40m] in the $X$ and $Y$ directions, and [-1m, 5.4m] along the $z$-axis, all within the ego coordinate system.
+This range is discretized into voxels, each with a size of [0.4m, 0.4m, 0.4m].
+Each occupied voxel is assigned one of 17 semantic labels, including 16 common classes and a general object class labeled as ‘others’.
+The dataset includes RGB images captured by six surrounding cameras and LiDAR point clouds, enabling dense voxel-wise annotations for 3D scene understanding.
+Additionally, the dataset provides visibility masks for both LiDAR and camera modes, which can be utilized for training purposes.
+SurroundOcc [51] is an automatically generated occupancy prediction dataset that does not require extra human annotation.
+It is built upon the nuScenes [3] and generates dense occupancy ground truth by utilizing existing 3D detection and 3D semantic segmentation labels.
+The dataset uses the nuScenes training set to train the model and employs the nuScenes validation set for evaluation purposes.
+The occupancy prediction range is set to [-50m, 50m] for the $X$ and $Y$ axes, and [-5m, 3m] for the $z$-axis, all within the LiDAR coordinate system.
+The final output 3D occupancy grid has a shape of 200×200×16, with each voxel measuring [0.5m, 0.5m, 0.5m].
+Evaluation metrics: Following previous methods, we use the mean Intersection over Union (mIoU) as our evaluation metric when training models with camera visible mask provided by Occ3D-nuScenes, ensuring a fair comparison with existing approaches.
+When the camera visible mask is excluded from the training phase, we report an additional semantic segmentation metric, RayIoU [31], which has been recently introduced.
+For the SurroundOcc dataset, we report semantic segmentation performance using IoU and mIoU.
+Here, IoU is defined as the intersection over union of occupied voxels, calculated while ignoring their semantic categories, and it is used as an evaluation metric for the scene completion (SC) task.
 4.2 Implementation Details
-We implement our method using PyTorch [36] and train all models on 8 NVIDIA RTX 4090 GPUs.
-We utilize ResNet-50 [11], which is pre-trained on nuImages [3], as the image backbone and crop the input image size to 256×704.
-In the LiDAR branch, we voxelize 10 LiDAR scans and start training from scratch.
-We employ the AdamW [32] optimizer with a cosine annealing learning rate scheduler, including a warmup phase, and set an initial learning rate of 2e-4.
-We apply data augmentation to both the input image and the ground truth 3D bounding boxes, and adopt CBGS [61].
-Figure 3: Qualitative visualizations on Occ3D-nuScenes validation set without camera mask.
+We use ResNet-50 [12], pre-trained on nuImages [3] by MMDetection [5], as our image backbone and adjust the input image resolution to 256×704.
+In the LiDAR branch, we aggregate LiDAR sweeps from 10 adjacent timestamps as input and initiate training from scratch.
+Data augmentation techniques, such as scaling, rotation, and flipping, are applied to the input images, and we also perform random flipping in the BEV space.
+The loss weights are configured as $\lambda_{l}=0.25$ and $\lambda=0.01$.
+We adopt the AdamW optimizer [35] with a cosine annealing learning rate scheduler, incorporating a warmup phase, and set the initial learning rate to 2e-4.
+Unless conducting ablation experiments, we train our DAOcc using CBGS [65] for 6 epochs, the input point cloud range is set to [-54.0, -54.0, -5.0, 54.0, 54.0, 3.0] and the voxel size is set to [0.075, 0.075, 0.2].
+All models are trained on 8 NVIDIA RTX4090 GPUs with a batch size of 4.
+Additionally, if the camera visible mask is not used during training, we randomly discard 80% of empty voxels, consistent with Zhao et al.
+[63].
+Temporal information fusion, exponential moving average (EMA) for weight parameter and test-time augmentation is not used.
 4.3 Comparison with State-of-the-Art Methods
+Method Modality 3D Backbone Param mIoU
+FB-Occ [24] C - 130.8M 48.90
+RadOcc [60] C+L SECOND 135.4M 49.38
+EFFOcc [43] C+L SECOND 108.6M 52.62
+DAOcc C+L SECOND 56.2M 53.82
+Table 4: Complexity comparisons.
+The C and L denote camera, LiDAR, respectively.
+3D Backbone represents the network used for point cloud feature extraction.
+The reported performance mIoU is the result of training with camera mask on Occ3D-nuScenes validation set.
 Comparison on Occ3D-nuScenes with camera mask.
-As presented in Table 1, we compare our DAOcc with both single-modal and multi-modal methods on Occ3D-nuScenes, all results are either implemented directly by their authors or reported based on their official code implementations.
-All methods use the camera visible mask during training, except for CTF-Occ [41].
-Our DAOcc establishes a new state-of-the-art performance, achieving an outstanding 53.82 mIoU, while utilizing a 2D backbone ResNet50 and 256×704 input image resolution.
-Compared with the current state-of-the-art multi-modal method EFFOcc [40], our DAOcc outperforms EFFOcc by 1.2 mIoU.
-Moreover, while EFFOcc employs Swin-B and an input image resolution of 512×1408, DAOcc utilizes only ResNet50 and an input image resolution of 256×704.
-In addition, EFFOcc requires the LiDAR branch to perform 3D object detection pre-training, which results in an improvement of approximately 3 mIoU, in order to achieve optimal results, whereas DAOcc does not require 3D detection pretraining.
+As presented in Table 1, we compare our proposed DAOcc with both image-based and multi-modal methods on the Occ3D-nuScenes dataset.
+All results are provided either directly by their authors or based on their official code implementations.
+Our DAOcc establishes a new state-of-the-art performance, achieving an outstanding 53.82 mIoU, while utilizing a 2D backbone ResNet50 [12] and 256×704 input image resolution.
+Compared with the latest state-of-the-art multi-modal method EFFOcc [43], DAOcc achieves improvement of 1.2 mIoU.
+Notably, the image feature extraction network Swin-B [32] in EFFOcc has approximately 3 times the number of parameters of ResNet50, and EFFOcc uses a larger input image resolution of 512×1408.
+This highlights our method’s ability to mine more useful information from point cloud data.
+In addition, to achieve optimal results, EFFOcc requires the LiDAR branch to undergo 3D object detection pre-training, which leads to an improvement of approximately 3 mIoU.
+In contrast, DAOcc does not require LiDAR branch pre-training, but instead seamlessly integrates 3D object detection as an auxiliary branch into the multi-modal occupancy prediction framework, simplifying the model training process and achieving better performance.
+Furthermore, we compare the complexity of the models, as shown in Table 4.
+Compared to the two top-performing multi-modal methods, RadOcc [60] and EFFOcc [43], as well as the top-performing image-based method FB-Occ [24], our proposed DAOcc has only about half the number of parameters but achieves significantly better performance (outperforming RadOcc, EFFOcc, and FB-Occ by 4.44, 1.2 and 4.92 mIoU).
 Comparison on Occ3D-nuScenes without camera mask.
 To the best of our knowledge, no existing multi-modal occupancy prediction work has used RayIoU metrics to report model performance, so we mainly compare with image-based methods.
-As shown in Table 2, thanks to the multi-modal input and the effectiveness of our proposed method, our DAOcc is significantly ahead of image-based methods, although many of these image-based methods use multi-frame temporal fusion.
-It should be noted that some image-based methods based on monocular depth estimation, such as Panoptic-FlashOcc [55], only require images as input during the inference phase, but require additional point cloud data input to provide depth information supervision during the training phase, so these methods can also be regarded as multi-modal during the training phase.
+As shown in Table 2, benefiting from the effective use of multi-modal input, especially point cloud data, our DAOcc achieves an outstanding 48.2 RayIoU and significantly outperforms image-based methods by a large margin, even though many of these methods use multi-frame temporal fusion.
+It should be noted that some image-based methods, such as Panoptic-FlashOcc and FB-Occ [24], adopt the Explicit Depth Supervision [22] and leverage depth information derived from point cloud data to supervise the training of monocular depth estimation from images, so these methods can also be regarded as multi-modal during the training phase.
 Comparison on SurroundOcc.
-Table 3 presents a quantitative comparison on SurroundOcc validation set, showcasing the performance of our DAOcc against other methods.
-By leveraging both LiDAR and camera inputs, DAOcc achieves a remarkable 3.2 mIoU increase over OccFusion [33], which not only employs camera and LiDAR but also incorporates radars.
-Notably, OccFusion utilizes R101-DCN with a high-resolution input of 900×1600, whereas our method employs a more lightweight Resnet50 backbone and a lower resolution of 256×704.
+Table 3 presents a quantitative comparison on the SurroundOcc validation set, showcasing the performance of our DAOcc against other methods.
+Leveraging both camera and LiDAR inputs, DAOcc achieves a notable 3.2 mIoU improvement over OccFusion [37], which incorporates radar data in addition to camera and LiDAR inputs.
+It is worth noting that OccFusion utilizes R101-DCN with a high-resolution input of 900×1600, whereas our method employs a more lightweight Resnet50 backbone and a lower input resolution of 256×704.
 This underscores the effectiveness and efficiency of our proposed method.
 4.4 Ablation Studies
-Ablation studies are conducted on Occ3D-nuScenes using the camera mask. Unless otherwise specified, all experiments presented in this section are trained for 15 epochs only.
-BEV View Range Voxel Resolution (BEV) mIoU
+Ablation studies are conducted on the validation set of Occ3D-nuScenes, with the camera visible mask being used during training.
+Unless otherwise specified, all experiments presented in this section are trained for 15 epochs without employing CBGS [65].
+BEV View Range (m) BEV Voxel Size (m) mIoU
 (a) $\left[-41.4,41.4\right]$ 0.075 $\times$ 0.075 50.76
-(b) $\left[-51.2,51.2\right]$ 0.1 $\times$ 0.1 49.98
-(c) $\left[-54.0,54.0\right]$ 0.075 $\times$ 0.075 51.26
-Table 4: Performance comparison between different BEV processing ranges and voxel sizes.
-We propose the BVRE strategy to enrich the spatial contextual semantic information by broadening the perceptual range from the BEV perspective.
-As shown in Table 4, comparing (a) and (c), increasing the BEV view range from [-41.4, 41.4] to [-54.0, 54.0] can bring a performance improvement of 0.5 mIoU.
-In addition, by comparing (a) and (b), we can find that the size of the voxel has a great impact on the model’s performance.
-Setting a smaller voxel size will bring obvious performance improvements but will also impose a greater computational burden.
-Detection Auxiliary mIoU bicycle motorcycle traffic cone
-51.26 20.60 38.89 36.65
-52.82 36.35 45.48 41.98
-Table 5: Ablation study on the auxiliary detection head.
-To fully leverage the inherent geometric structures within point cloud features, we incorporate 3D object detection as auxiliary supervision.
-As shown in Table 5, by adding auxiliary supervision for 3D object detection, we can achieve an improvement of 1.56 mIoU, demonstrating the effectiveness of our method.
-Given that the gain in object detection for foreground objects will be greater, we present the three categories with the highest gain in foreground objects, which are the bicycle (+15.75 mIoU), the motorcycle (+6.59 mIoU) and the traffic cone (+5.33 mIoU).
-This further proves the rationality of the performance improvement of our method.
-Detection Auxiliary EP15 EP24 CBGS (EP6) mIoU
+(b) $\left[-45.6,45.6\right]$ 0.075 $\times$ 0.075 50.98
+(c) $\left[-45.6,45.6\right]$ 0.1 $\times$ 0.1 49.75
+(d) $\left[-51.2,51.2\right]$ 0.1 $\times$ 0.1 49.98
+(e) $\left[-54.0,54.0\right]$ 0.075 $\times$ 0.075 51.26
+Table 5: Performance comparisons between different BEV ranges and voxel sizes. In all experiments, the voxel height is 0.2m.
+Effectiveness of the BVRE.
+We propose the BVRE strategy to enrich spatial contextual information by expanding the perceptual range from a BEV (Bird’s Eye View) perspective.
+As shown in Table 5, a comparison of (a) and (b) indicates that, with the same voxel size, expanding the BEV range from [-41.4m, 41.4m] to [-45.6m, 45.6m] yields a performance improvement of 0.22 mIoU.
+Furthermore, the comparison of (a) and (e) shows that further expanding the BEV range to [-54.0m, 54.0m] augments the performance gain to 0.5 mIoU.
+This observation validates the value of spatial context information in a larger BEV range.
+In addition, by comparing (b) and (c), we can find that the size of the voxel has a great impact on the model’s performance. To achieve better performance, we adopt the setting (e) in DAOcc.
+camera mask ADH mIoU
+$\blacksquare$ bicycle
+$\blacksquare$ motorcycle
+$\blacksquare$ pedestrian
+$\blacksquare$ traffic cone
+51.26 20.60 38.89 55.44 36.65
+52.82 (+1.56) 36.35 (+15.75) 45.48 (+6.59) 58.27 (+2.83) 41.98 (+5.33)
+44.31 14.94 32.22 43.77 34.07
+46.89 (+2.58) 32.51 (+17.57) 42.89 (+10.67) 53.41 (+9.64) 40.55 (+6.48)
+Table 6: Ablation study on the Auxiliary Detection Head. ADH is short for the Auxiliary Detection Head.
+Ablation on the Auxiliary Detection Head.
+To fully leverage the inherent geometry and structure information within point cloud features, we incorporate 3D object detection as auxiliary supervision.
+As shown in Table 6, by adding auxiliary 3D object detection supervision, we can achieve an improvement of 1.56 mIoU when using camera visible masks during training and 2.58 mIoU when not using the mask, respectively.
+This demonstrates that adding a detection branch can significantly enhance the discriminability of the multi-modal features.
+Furthermore, we report on the four categories that benefit the most from the 3D object detection supervision.
+Notably, all four categories are foreground objects, which further validates the effectiveness and rationale behind the performance improvement achieved by our method.
+Specifically, when the camera visible mask is used during training, the addition of the auxiliary detection branch results in improvements of 15.75 mIoU, 6.59 mIoU, 2.83 mIoU, and 5.33 mIoU for bicycles, motorcycles, pedestrians, and traffic cones, respectively.
+Conversely, when the camera visible mask is not used during training, the improvements are 17.57 mIoU, 10.67 mIoU, 9.64 mIoU, and 6.48 mIoU for the same categories, respectively.
+5 Conclusion
+In this paper, we propose a novel multi-modal occupancy prediction framework, DAOcc, which aims to achieve superior performance while utilizing a deployment-friendly image backbone and practical input image resolutions.
+DAOcc achieves this superior performance by fully leveraging the unique advantages of point cloud data within a multi-modal framework, advantages that have been largely overlooked by previous methods.
+Extensive experiments conducted on the Occ3D-nuScenes and SurroundOcc datasets demonstrate the superiority of DAOcc compared to existing methods.
+We believe that DAOcc has practical application potential and could inspire future research in multi-modal occupancy prediction.
+Supplementary Material
+Method Modality GPU Type FPS mIoU
+FlashOcc(2f) [57] C A100-PCIE-40G 0.9 43.52
+RadOcc [60] C+L - 0.3 49.38
+EFFOcc [43] C+L RTX A6000 1.8 52.62
+DAOcc(ours) C+L RTX 3070Ti 2.7 53.82
+Table 7: Comparisons of the runtime.
+(xf) means use x frames for temporal fusion.
+The C and L denote camera, LiDAR, respectively.
+FPS for DAOcc is measured with the PyTorch fp32 backend.
+The reported performance mIoU is the result of training with camera mask on Occ3D-nuScenes validation set.
+6 Runtime Comparisons
+We compare the runtime of different methods, as shown in Table 7.
+The FPS is obtained by testing on the corresponding GPU Type listed in the table.
+Since the source code for RadOcc is unavailable and EFFOcc has not yet released its implementation, we rely on the runtime speeds reported in their respective papers as benchmarks.
+It’s worth noting that the computing capability of the RTX 3070 Ti is significantly lower than that of both the RTX A6000 and the A100.
+The results demonstrate that our proposed DAOcc achieves the best performance while maintaining the fastest runtime speed.
+7 Ablation for Achieving Final Performance
+In addition to the ablation studies presented in Section 4.4, this section provides further ablation experiments.
+Table 8 presents the ablation experiments that contributed to the final performance.
+First, by extending the training time from 15 epochs to 24 epochs, we can observe an improvement of 0.57 mIoU.
+Next, the use of CBGS [65] yields an additional 0.43 mIoU improvement.
+Since CBGS involves resampling, which increases the number of samples per epoch, we limit the training to 6 epochs to maintain a comparable total training time to the 24 epochs without CBGS for a fair comparison.
+8 Detection Performance
+Since our proposed DAOcc incorporates an auxiliary 3D object detection branch into the occupancy prediction framework, we also evaluate the performance of this auxiliary branch.
+The auxiliary detection branch is designed to identify 10 object categories, including cars, trucks, buses, trailers, construction vehicles, pedestrians, motorcycles, bicycles, traffic cones, and barriers.
+All experiments are conducted using the nuScenes [3] dataset for the auxiliary 3D object detection training, and the results are evaluated using the mean average precision (mAP) and nuScenes detection score (NDS) metrics provided by the nuScenes benchmark.
+As shown in Table 9, when DAOcc is trained on the Occ3D-nuScenes [44] dataset (for occupancy prediction) and the camera visible mask is used during training, the auxiliary detection branch achieves a performance of 59.4 mAP and 64.3 NDS.
+Notably, comparable performance is observed when the camera visible mask is not used.
+However, when the SurroundOcc [51] dataset is employed for occupancy prediction training, the auxiliary detection branch achieves a lower performance of 53.6 mAP and 60.2 NDS.
+We attribute this performance gap to the fact that the surroundocc dataset is automatically generated without manual calibration, potentially introducing more noise and negatively affecting model convergence.
+EP15 EP24 CBGS (EP6) mIoU
 52.82
@@
 53.82
-Table 6: Ablation for obtaining the final performance.
-Table 6 provides ablation experiments to obtain the final performance.
-First, by increasing the training time from 15 epochs to 24 epochs, an improvement of 0.57 mIoU can be achieved.
-Then, by using CBGS [61], another improvement of 0.43 mIoU can be achieved.
-When using CBGS, since resampling increases the number of samples in each epoch, for the sake of a fair comparison, we limited the training to 6 epochs to approximate the total training time of 24 epochs without CBGS.
-4.5 Visualization
-Figure 3 presents the visualizations of DAOcc on Occ3D-nuScenes validation set without using camera mask.
-The results show that our DAOcc made a relatively complete prediction of the scene, accurately reproducing it in fine detail.
-5 Conclusion
-In this paper, we propose a novel multi-modal occupancy prediction framework, aiming to achieve superior performance while using a deployment-friendly image feature extraction network and input image resolution.
-Extensive experiments on Occ3D-nuScenes and SurroundOcc demonstrate the superiority of our approach over existing methods.
-We believe that DAOcc can be put into practice and inspire future research.
+Table 8: Ablation for achieving final performance. EP15, EP24, and EP6 represent training for 15 epochs, 24 epochs, and 6 epochs, respectively.
+9 Visualization
+Figure 4 presents the visualizations of DAOcc on Occ3D-nuScenes validation set.
+The camera visible mask is not used during model training and visualization.
+The results demonstrate that DAOcc is capable of making a relatively complete prediction of the scene, accurately reproducing it in fine detail.
+Dataset Camera Mask mIoU mAP NDS Car Truck Bus T.L. C.V. Ped. M.C. B.C. T.C. B.R.
+Occ3D-nuScenes [44] 53.8 59.4 64.3 85.7 57.2 65.8 33.1 22.8 75.5 68.3 47.3 70.2 68.2
+Occ3D-nuScenes [44] 47.9 59.6 64.2 85.4 55.0 64.6 34.1 23.1 76.3 67.0 50.7 70.8 69.0
+SurroundOcc [51] - 30.5 53.6 60.2 81.9 42.7 56.0 26.3 19.7 72.6 57.7 43.8 66.7 68.1
+Table 9: 3D occupancy prediction and detection performance of DAOcc.
+‘T.L.’, ‘C.V.’, ‘Ped.’, ‘M.C.’, ‘B.C.’, ‘T.C.’, and ‘B.R.’ are short for trailer, construction vehicle, pedestrian, motorcycle, bicycle, traffic cone and barrier, respectively.
+Figure 4: Qualitative visualizations on Occ3D-nuScenes validation set. The camera visible mask is not used during model training and visualization.